# AV-Link: Temporally-Aligned Diffusion Features for Cross-Modal Audio-Video Generation

### v1

@@ Abstract
We propose AV-Link, a unified framework for Video-to-Audio and Audio-to-Video generation that leverages the activations of frozen video and audio diffusion models for temporally-aligned cross-modal conditioning. The key to our framework is a Fusion Block that enables bidirectional information exchange between our backbone video and audio diffusion models through a temporally-aligned self attention operation. Unlike prior work that uses feature extractors pretrained for other tasks for the conditioning signal, AV-Link can directly leverage features obtained by the complementary modality in a single framework i.e. video features to generate audio, or audio features to generate video. We extensively evaluate our design choices and demonstrate the ability of our method to achieve synchronized and high-quality audiovisual content, showcasing its potential for applications in immersive media generation.
@@ 1 Introduction
Generative models are becoming proficient at synthesizing high-quality video [6, 78, 58, 70, 53, 33] and audio [23, 19, 48, 49, 79], producing content that is increasingly realistic and contextually coherent. As a result, recent work has tackled the challenging task of Video-to-Audio (V2A) generation where the goal is to add sound that is both semantically and temporally aligned to video content [52, 97, 85, 62]. Current V2A models [52, 97, 85, 62] typically rely on pretrained visual feature extractors such as CLIP [63], MetaCLIP [86], ImageBind [21], and CAVP [52] to condition audio generation and create soundscapes that are semantically aligned with the visual frames.
An even more challenging task is Audio-to-Video (A2V) where the goal is to generate video that could plausibly visualize the sounds in a given input audio. Although this task is even more ambiguous, some recent work has demonstrated encouraging results [93, 96, 41]. similar to V2A, these methods have explored conditioning video generators on audio features extracted with pretrained audio models such as ImageBind [21], CLAP [16] and BEATs [10]. While the use of pretrained feature extractors has been important in achieving semantic alignment, these features do not exhibit precise temporal alignment capabilities, a fundamental aspect of cross-modal audio-video generation. Moreover, since these models require feature extractors pre-trained on complementary data for each modality, it forces a separate treatment for V2A and A2V tasks.
To address these limitations, recent work generates audio and video jointly within a single model [67, 35] to unify the multimodal generative task. Although promising in concept, these models have faced considerable challenges and present a performance gap with respect to V2A and A2V which we show can be reconducted to the absence of informative multimodal features in early diffusion timesteps.
@@
In this work, we tackle cross-modal audiovisual generation within a unified architecture named AV-Link that is capable of semantically and temporally aligned generation for either V2A or A2V as shown in Fig. 1. Starting from two well-trained flow models for audio and video generation, we use their activations as conditioning signals for either V2A or A2V synthesis, bypassing the need for pretrained feature extractors. Such activations contain rich semantics [51], but most importantly they are also temporally-aligned as generated media (e.g. audio and video) originates from such activations. This approach is supported by prior findings which showed that diffusion activations are semantically rich spatially aligned, making them useful for downstream pixel-aligned tasks such as semantic keypoint matching [51, 26] segmentation [91] and depth estimation [34].
To make use of such features, we freeze the pretrained generators and introduce a Fusion Block that enables bidirectional information exchange between audio and video modalities by connecting activations from the pretrained generators through a temporally-aligned self-attention operation and conditioning both backbones through symmetric feature reinjection (see Fig. 2). This approach offers superior semantic and temporal alignment compared to pre-trained features alone. Another benefit is that it fully leverages the capabilities of the pretrained audio and video networks and only adds a set of Fusion Blocks, which is relatively small. This compact design simplifies deployment for applications requiring comprehensive functionality: video, audio, video-to-audio, and audio-to-video generation.
To foster future work, we present insights into diffusion features and cross-modal architectures. First, we inspect activations to identify noise levels and model blocks that yield the most effective conditioning features. Next, we explore different mechanisms for injecting such features into the generators. Finally, we show that diffusion features outperform specialized pretrained feature extractors.
@@
In the V2A task, when evaluated against the state-of-the-art on VGGSounds [9] our method shows the best overall performance and drastically improves temporal alignment as measured by Onset ACC [15] by up to $36.5\%$ points over the best baseline. User studies show a strong preference for our method which is preferred over Movie Gen Audio [62] $56.8\%$ of the times with respect to semantic alignment, and $63.6\%$ with regards to temporal alignment. In the V2A task, our method strongly surpasses TempoTokens [77] in both video quality and audio-video alignment, with user studies showing an overwhelming preference for our method.
To summarize, we introduce the following contributions:
  • AV-Link, a unified framework to address Audio-to-Video and Video-to-Audio tasks. For the first time in this context, we propose the use of diffusion features to sidestep the need for specialized pretrained feature extractors and to improve temporal alignment.
  • A Fusion Block based on a temporally-aligned self attention operation and symmetric feature reinjection mechanism to allow bidirectional information exchange between the frozen generators for each modality
  • To foster future work in this direction, we present extensive ablations into diffusion feature selection strategies and mechanisms for cross-modal feature injection and show that such features outperform commonly used pretrained feature extractors.
@@ 2 Related Work
Text-to-Audio Generation Text-to-audio generation methods [54, 88, 19, 49, 29, 23] aim to produce audio signals based on a given text condition. Common approaches [29, 23] leverage latent diffusion models [25], where they represent audio as Mel-spectrograms, encoding them with a 2D Variational Autoencoder (VAE) and using a U-Net architecture for denoising. Successive works employ scalable transformer-based diffusion backbones to achieve superior quality. Make-an-Audio 2 [29] encodes the spectrogram with a 1D VAE and uses a transformer backbone. GenAU [23] introduced a scalable dataset collection framework and transformer architecture to produce high-quality sound effects. Movie Gen [62] directly encodes waveforms with a VAE, bypassing intermediate Mel-spectrogram representations to enable generation at a higher resolution of 44kHz. Such approaches operate in a text-conditioned setting [19, 49, 29], and showed the importance of text-encoding in achieving better quality and text-audio alignment. Our work focuses on establishing a unified framework for Video-to-Audio and Audio-to-Video leveraging diffusion model activations as the main conditioning signal, posing a more complex challenge due to the need for precise temporal alignment with the conditioning modality.
@@
Video-to-Audio Generation The task of generating audio from silent videos recently emerged as a difficult generative task [81, 12, 59, 57, 73, 40, 56, 65, 2, 89, 8, 7, 22, 11, 38, 46, 44, 94, 74] and gained relevance not only for its potential in simplifying movie production but also for the recent advent of high-quality video generation methods [58, 62, 92, 33, 43, 99, 6, 3, 4] which produce no audio output. Video feature representations play a crucial role in achieving highly aligned generation and have been the focus of most recent work [82, 31, 69, 84, 80, 52, 97, 62, 66, 30]. SpecVQGAN [31] employs a cross-modal transformer conditioned on ImageNet-pretrained visual backbones. Im2Wav [69] uses per-frame CLIP [63] features to condition audio generation. SonicVisionLM [84] bridges the visual and audio modalities through a pretrained textual representation. V2A-Mapper [80] aligns the CLIP [63] and CLAP [16] to leverage a pretrained text-to-audio model for V2A. Diff-Foley [52] and Frieren [82] address alignment using CAVP, an ad-hoc contrastive video-audio representation. FoleyCrafter [97] augments a frozen audio generator with a learnable semantic adapter and temporal controller. Movie Gen Audio [62] obtains strong video-audio alignment by conditioning on a descriptive text caption and per-frame MetaCLIP [86] features. Despite such advancements, obtaining precise temporal alignment remains challenging. We show that activations extracted from diffusion backbones provide rich and temporally aligned representations that sidestep the need for ad-hoc extractors.
@@
Audio-to-Visual Generation Similar to the Video-to-Audio task, recent work has focused on building suitable audio representations, paying special attention to time alignment. Lee et al. [42] employ a Sound Inversion Encoder to map audio features to the StyleGAN latent space. Seeing-and-Hearing [85] propose a video-audio aligner based on ImageBind [21]. TPoS [32] manipulates generated images conditioned on audio representations extracted through an ad-hoc temporally-aligned audio encoder. TempoTokens [93] aligns a frozen video generator to the audio modality by employing features from a finetuned BEATs encoder. AADif [41] leverages the CLAP [16] audio encoder and audio magnitude to condition generation. Recently, AVSyncD [96] encodes audio into temporally-dependent tokens using ImageBind [21]. We show that temporally-aligned Video-to-Audio and Audio-to-Video generation can be achieved within the same framework by leveraging existing diffusion activations without the need for specialized feature representations.
@@
Joint Video-Audio Generation Recently, several works emerged that generate the audio and video modalities jointly. [24, 36, 90, 45, 75, 72] MM-Diffusion [67] proposes a diffusion framework based on a Coupled U-Net for simultaneously denoising the audio and video streams. Seeing-and-Hearing [85] employs ImageBind [21] to establish alignment between modalities. AVDiT [35] introduces a multimodal DiT design and learns multiple conditional distributions over audio and video through a mixture of noise levels, which enables the framework to perform a range of tasks including joint audio-video generation, Video-to-Audio, and Audio-to-Video. Such approaches, however present lower Video-to-Audio performance than ad-hoc methods [97, 62], thus in this work, we focus on the conditional generation tasks.
@@ 3 Method
Figure 2: Design of the proposed Fusion Block connecting the frozen video and audio backbones. A RoPE-based temporal alignment mechanism establishes correspondences between modalities that are leveraged by self attention. Video and audio features are symmetrically reinjected into the frozen generators. The block is regularly applied multiple times throughout the backbones.
Starting with pretrained and frozen Text-to-Audio (T2A) and Text-to-Video (T2V) generators, we aim to combine the two under a unified architecture to produce either Video-to-Audio (V2A) $\mathcal{G}_{V2A}$ or Audio-to-Video (A2V) $\mathcal{G}_{A2V}$ generators. To achieve this, we propose AV-Link, a symmetric architecture that benefits from the activations of one generator to condition the generator of the other modality. In Sec. 3.1, we describe the background on Flow Matching models. We then describe in Sec. 3.2, the architecture of our pretrained audio and video generators. Finally, in Sec. 3.3, we describe the architecture of the multimodal Fusion Blocks that connect the two generators.
@@ 3.1 Background
where $p_{t}$ indicates a training distribution over time $t$, which we instantiate as a logit normal distribution [17].
At inference time, an ODE solver such as first-order Euler can be employed to produce samples $\mathbf{X}_{1}$ starting from Gaussian noise $\mathbf{X}_{0}$ using the model’s velocity estimates.
@@ 3.2 Base Models
Audio Model. Given an audio signal from audio dataset $\mathcal{D}_{A}$, we produce a corresponding Mel-spectrogram and adopt a 1D-VAE [23] to encode it into a 1D sequence of latent. The latent audio representation modeling $\mathbf{A}_{1}$ thus has shape $\mathbb{R}^{T^{l}_{a}\times D^{l}_{a}}$, where $T^{l}_{a}$ and $D^{l}_{a}$ respectively indicate the number of tokens and their dimensionality produced by the autoencoder. The sequence is modeled by a DiT [60] (see Fig. 2) consisting of a stack of identical blocks containing a self attention operation followed by an MLP. Information on the current timestep $t$ is injected through adaptive layernorm [61] and the attention operator is augmented with 1D-RoPE [71] to encode positional information. To condition on text prompts, we encode them using the T5 [64] encoder and insert a cross attention layer attending to the text embeddings after each self attention operation [62]. The model contains 576M parameters and is trained to generate variable-length audio clips at 16kHz. We train it using the Flow Matching objective (Eq. (2)) over 100k iterations, with a learning rate of $43e\!-\!$ 4 and a batch size of 1024, distributed across 8 A100 GPUs. During inference, a pre-trained Vocoder [37] converts the generated Mel-spectrograms into waveforms.
Video Model. Given an RGB video $\mathbf{V}_{1}$, we flatten it to shape $\mathbb{R}^{T^{l}_{v}\times 3}$, where $T^{l}_{v}$ indicates the total number of video pixels. We adopt a DiT architecture symmetric to the one of the audio model, where 3D-RoPE embeddings [13] encode positional information. Full self attention is employed to capture spatial and temporal relationships. The model has 576M parameters and is trained to generate 5.16-second videos at a resolution of 36 x 64 pixels at 6 frames-per-seconds, using the Flow Matching objective in Eq. (2). Training is conducted for 250,000 iterations with a learning rate of 0.0003 and a batch size of 512 on 16 A100 GPUs.
@@ 3.3 Multimodal Fusion Block
Cross-modal generation requires a high-quality representation for the conditioning modality [52, 97]. In the context of V2A, obtaining high-quality video features has been the focus of recent works, which explored the usage of contrastive representations such as ImageBind [85], CLIP [97], MetaCLIP [62], or built ad hoc video representations [52, 97]. While such representations capture the overall video semantic, we observe that they lack precise audio-video temporal alignment, producing characteristic temporal misalignment (see Sec. 4.4). On the other hand, the video generator is capable of generating videos from scratch, implying that its activations contain semantically and temporally aligned information that can be leveraged for V2A.
We thus propose a fully symmetric framework for V2A and A2V that leverages activations from the frozen generator of the conditioning modality to achieve aligned cross-modal generation using no ad-hoc feature extractors.
Fusion Blocks We address V2A and A2V by linking frozen audio and video generators $\mathcal{G}_{A}$, $\mathcal{G}_{V}$ through the Fusion Block depicted in Fig. 2. Consider the output activations of audio and video DiT blocks $\boldsymbol{x}_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\boldsymbol{x}_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$ at diffusion timestep $t_{a}$, $t_{v}$, respectively and notice one that one of the two, depending on the task, is always derived from ground truth conditioning input. The Fusion Block produces video-aware audio activations $\boldsymbol{\hat{x}}_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ and audio-aware video activations $\boldsymbol{\hat{x}}_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$:
@@
The Fusion Block presents a symmetric design that equally treats the different modalities. Initially, we project activations into a common dimension $D$ to produce $\bar{\boldsymbol{x}}_{a}\in\mathbb{R}^{T_{a}\times D}$, $\bar{\boldsymbol{x}}_{v}\in\mathbb{R}^{T_{v}\times D}$ respectively. Then, a multi-head self attention block is applied to the concatenation of audio and video features $\{\bar{\boldsymbol{x}}_{a},\bar{\boldsymbol{x}}_{v}\}$. Finally, an MLP projects the output back to the original dimensionality. Different sets of learnable parameters are used for tokens of each modality to allow for better modality alignment [17].
Symmetric Feature Reinjection Both $\boldsymbol{\hat{x}}_{a}$ and $\boldsymbol{\hat{x}}_{v}$ are passed to the subsequent DiT blocks in the respective frozen backbone to inject cross-modal information. Such design, which we call symmetric feature reinjection, allows us to further improve the quality of the activations extracted from successive blocks from the conditioning modality. In contrast to using static feature extractors [62, 52, 97], symmetric feature reinjection continuously refines the conditioning features according to the generated modality.
Multimodal Temporal Alignment To facilitate temporal alignment between modalities, we propose the use of temporally-aligned 1D RoPE [71] embeddings in the self attention operation. We express 1D temporal RoPE as:
@@
Where $\eta_{a}$, $\eta_{v}$ are the number of tokens to represent one second of audio and video respectively. Intuitively, the time-aligned embeddings ensure that each temporally corresponding audio and video tokens are rotated by RoPE by the same angle, set to be proportional to their position on the temporal axis, thus establishing temporal correspondence which facilitates temporally-synced multimodal information exchange.
Time-Aware Feature Fusion Information on the flow timesteps for audio $t_{a}$ and video $t_{v}$ are injected through adaptive adaLN-Zero [60], with audio and video parameter sets receiving information on both timesteps. Such design creates conditioning features that are most suited for the current flow time for the generated modality, increasing the effectiveness of symmetric feature reinjection.
@@
Inference In the V2A setting, we consider the input conditioning video $\mathbf{V}_{1}$ and compute partially noised video $\mathbf{V}_{t_{v}}$ at the chosen flow time $t_{v}$, which we set to a fixed level during the entire inference phase to serve as the input for the video backbone. As we later show in 4.4, it is crucial to inject some noise into the conditioning modality for extracting useful features from the conditioning model. We sample $\mathbf{A}_{0}$ from Gaussian noise at $t_{a}\!=\!0$ and use velocity predictions from $\mathcal{G}_{V2A}$ to progressively denoise it into $\mathbf{A}_{1}$. We perform sampling using a first-order Euler solver with 64 sampling steps. Text prompts can optionally be used for both modalities to condition generation. We perform A2V inference completely symmetrically.

### v2

@@ Abstract
We propose AV-Link, a unified framework for Video-to-Audio (A2V) and Audio-to-Video (A2V) generation that leverages the activations of frozen video and audio diffusion models for temporally-aligned cross-modal conditioning. The key to our framework is a Fusion Block that facilitates bidirectional information exchange between video and audio diffusion models through temporally-aligned self attention operations. Unlike prior work that uses dedicated models for A2V and V2A tasks and relies on pretrained feature extractors, AV-Link achieves both tasks in a single self-contained framework, directly leveraging features obtained by the complementary modality (i.e. video features to generate audio, or audio features to generate video). Extensive automatic and subjective evaluations demonstrate that our method achieves a substantial improvement in audio-video synchronization, outperforming more expensive baselines such as the MovieGen video-to-audio model.
@@ 1 Introduction
Generative models are becoming proficient at synthesizing high-quality video [7, 84, 63, 76, 58, 38, 1] and audio [27, 23, 53, 54, 85]. Given the intrinsic correlation between audio and video, there has been a growing interest in cross-modal generation including Video-to-Audio (V2A) and Audio-to-Video (A2V). The goal of V2A models is to add sound that is both semantically and temporally aligned to video content [57, 104, 92, 68]. Conversely, A2V models [100, 103, 45] seek to generate video that plausibly represent the sounds of a given audio. Since text has been established as a robust conditioning modality to guide the semantics of generated audio and video [99, 32], achieving accurate audio–video temporal alignment remains as the most critical and challenging aspect in cross-modal generation. Current V2A models [57, 104, 92, 68] typically depend on pre-trained visual feature extractors such as CLIP [69], MetaCLIP [93], ImageBind [25], and CAVP [57] to condition audio generation and create soundscapes that are semantically aligned with visual frames. Similarly, A2V methods condition video generators on features from pretrained audio models such as ImageBind [25], and BEATs [11]. However, the reliance on modality-specific feature encoders necessitates distinct models for V2A and A2V, limiting the potential for a unified framework and complicating the deployment for applications requiring comprehensive capabilities: text-to-audio (T2A), text-to-video (T2V), V2A, and A2V.
Recognizing the advantages of a single framework capable of generating both audio and video, recent work explores training a single model for joint audio-video generation [73, 87, 40, 36]. Although promising, these models have a performance gap with respect to conditional models (i.e. dedicated V2A and A2V models), and often operate on simpler datasets such as landscape videos [73, 87, 46] or exhibit weaker audio-video synchronization [36].
@@
In this work, we introduce AV-Link, a unified framework that is capable of semantically and temporally aligned generation for both V2A and A2V as shown in Fig. 1. Starting from two well-trained and frozen flow models for audio and video generation, we propose to use their activations as conditioning signals for V2A or A2V synthesis, bypassing the need for pretrained feature extractors. Such activations contain rich semantics [56], but they are also temporally-aligned since these activations are trained to generate modalities with a temporal component (i.e. audio and video). This is motivated by prior findings showing that activations in image diffusion models encode substantial semantic and spatial information, making them useful for pixel-aligned tasks such as semantic keypoint matching [56, 29] segmentation [98] and depth estimation [39].
To preserve the quality of the frozen unimodal generators while enabling cross-modal generation, we introduce a Fusion Block that connects the audio and video modalities by conditioning the generation of one modality on the activations from the other. Furthermore, we propose time-aligned Rotary Position Embedding (RoPE) as a flexible mechanism to align audio and video tokens across the temporal axis, demonstrating its benefits over prior approaches for temporal alignment. Moreover, although static diffusion model activations serve as a robust conditioning signal, cross-modal generation may benefit from a conditioning signal that dynamically evolves throughout the generation process. To this end, we introduce a symmetric feature reinjection strategy that leverages the diffusion prior of the pretrained model for the conditioning modality to refine the conditioning signal, enabling bidirectional information exchange between the conditioning and generated modality. By introducing a relatively small set of parameters (186M parameters), AV-Link offers a compact design that simplifies training and deployment for applications requiring comprehensive T2V, T2A, V2A, and A2V generation. Rather than focusing on improving single-modality generation—a topic extensively explored in concurrent studies [27, 99]—this work builds on those efforts by unifying pretrained audio and video generators in a single framework, emphasizing temporal alignment as the core challenge in cross audio-visual generation.
@@
In summary, we propose AV-Link, a unified framework to address A2V and V2A tasks. We show for the first time that video and audio diffusion activations contain rich temporal information and can replace specialized pretrained feature extractors while improving temporal alignment. The key to our method is a novel design of a Fusion Block that enables bidirectional information exchange between the conditioning and generated modalities through a symmetric feature reinjection mechanism. When evaluated against the state-of-the-art V2A method on VGGSounds [10], our method shows the best overall performance and drastically improves temporal alignment as measured by Onset ACC [20] by up to $76.4\%$ over the best baseline. User studies show a strong preference for our method which is preferred over Movie Gen Audio [68] (13B parameters) $63.6\%$ of the time with regards to temporal alignment. In the A2V task, AV-Link surpasses TempoTokens [83] in video quality and audio-video alignment, with user studies showing an overwhelming preference for our method.
@@ 2 Related Work
Text-to-Audio Generation Text-to-audio methods generate audio conditioned a given text [59, 95, 23, 54, 32, 27]. Recent works employ scalable transformer-based diffusion backbones over a latent representation of Mel-spectrograms to achieve superior quality. Make-an-Audio 2 [32] encodes the spectrogram with a 1D VAE and uses a DiT [66] backbone. GenAU [27] introduced a scalable dataset collection framework and transformer architecture to produce high-quality sound effects. Movie Gen [68] directly encodes waveforms with a VAE, bypassing intermediate Mel-spectrogram representations. Such approaches operate in a text-conditioned setting [23, 54, 32], and showed the importance of text encoding in achieving better quality and text-audio alignment. Our work focuses on establishing a unified framework for text-conditioned V2A and A2V, posing a more complex challenge due to the need for precise temporal alignment with the conditioning modality.
@@
Video-to-Audio Generation The task of generating audio from silent videos recently emerged as a difficult generative task [88, 17, 65, 62, 79, 44, 61, 71, 3, 96, 9, 8, 26, 15, 42, 51, 49, 101, 80] and gained relevance not only for its potential in simplifying movie production but also because of the recent development of high-quality video generation methods which produce no audio output [63, 68, 99, 38, 47, 106, 7, 4, 5]. Video feature representations play a crucial role in achieving highly aligned generation and have been the focus of most recent work [89, 35, 75, 91, 86, 57, 104, 68, 72, 33]. SpecVQGAN [35] employs a cross-modal transformer conditioned on ImageNet-pretrained visual backbones. V2A-Mapper [86] aligns CLIP [69] and CLAP [21] to leverage a pretrained text-to-audio model for V2A. Diff-Foley [57] and Frieren [89] address alignment using CAVP [13], an ad-hoc contrastive video-audio representation. FoleyCrafter [104] augments a frozen audio generator with a learnable semantic adapter and temporal controller. Movie Gen Audio [68] obtains strong video-audio alignment by conditioning on a descriptive text caption and per-frame MetaCLIP [93] features. Concurrent work [16, 14] demonstrated remarkable capabilities by leveraging training on audio-only datasets. Despite such advancements, the unification of A2V and V2A in a single framework remains a challenge. We show that temporally-aligned V2A and A2V generation can be achieved within the same framework by leveraging existing diffusion activations without the need for specialized feature extractors.
@@
Audio-to-Video Generation Similar to the Video-to-Audio task, recent work has focused on building suitable audio representations, paying special attention to time alignment. Lee et al. [46] employ a Sound Inversion Encoder to map audio features to the StyleGAN latent space. Seeing-and-Hearing [92] propose a video-audio aligner based on ImageBind [25]. TPoS [37] manipulates generated images conditioned on audio representations extracted through an ad-hoc temporally-aligned audio encoder. TempoTokens [100] aligns a frozen video generator to the audio modality by employing features from a finetuned BEATs encoder. AADif [45] leverages the CLAP [21] audio encoder and audio magnitude to condition generation. Recently, AVSyncD [103] encodes audio into temporally-dependent tokens using ImageBind [25]. We show that using activations from a pretrained audio generator provides a stronger conditioning signal than external feature extractors, yielding improved temporal alignment.
@@
Joint Video-Audio Generation Several works emerged that generate the audio and video modalities jointly. [28, 40, 97, 50, 81, 78] MM-Diffusion [73] proposes a diffusion framework based on a Coupled U-Net for simultaneously denoising the audio and video streams. Seeing-and-Hearing [92] employs ImageBind [25] to establish alignment between modalities. AVDiT [87] introduces a multimodal DiT design and learns multiple conditional distributions over audio and video through a mixture of noise levels, enabling V2A, A2V, and joint audio-video generation. Such approaches, however, present lower performance than conditional methods [104, 68]. Additionally, they are trained from scratch on audio-video datasets. Since most large-scale audio and video datasets are curated around a single modality, the other modality may be missing, have insufficient quality, or unlabeled [12, 27], this limits the data scalability of such methods. Thus, in this work, we target cross-modal generation (i.e. V2A and A2V) rather than joint audio-video generation.
@@ 3 Method
Figure 2: Design of the proposed Fusion Block connecting the frozen video and audio backbones. A RoPE-based temporal alignment mechanism aligns the representation of the two modalities which are processed by self attention. Video and audio features are symmetrically reinjected into the frozen generators. The block is regularly applied multiple times throughout the backbones.
Starting with pretrained and frozen Text-to-Audio and Text-to-Video generators, we aim to combine the two under a unified architecture to produce Video-to-Audio $\mathcal{G}_{V2A}$ and Audio-to-Video $\mathcal{G}_{A2V}$ generators. To achieve this, we propose AV-Link, a symmetric architecture that benefits from the activations of one generator to condition the generator of the other modality. In Sec. 3.1, we describe the background on Flow Matching models. We then describe in Sec. 3.2, the architecture of our pretrained audio and video generators. Finally, in Sec. 3.3, we describe the architecture of the multimodal Fusion Blocks that connect the two generators.
@@ 3.1 Background
where $p_{t}$ indicates a training distribution over time $t$, which we instantiate as a logit normal distribution [22]. At inference time, an ODE solver such as a first-order Euler solver can be employed to produce samples $\mathbf{X}_{1}$ starting from Gaussian noise $\mathbf{X}_{0}$ using the model velocity estimates.
@@ 3.2 Base Models
Audio Model. Given an audio signal from audio dataset $\mathcal{D}_{A}$, we produce a corresponding Mel-spectrogram and adopt a 1D-VAE [27] to encode it into a 1D sequence of latent representations. This latent audio representation $\mathbf{A}_{1}$ thus has shape $\mathbb{R}^{T_{a}\times D_{a}}$, where $T_{a}$ and $D_{a}$ respectively indicate the number of tokens and their dimensionality produced by the autoencoder. The sequence is modeled by a DiT [66] (see Fig. 2) consisting of a stack of identical blocks containing a self attention operation followed by an MLP. Information on the current timestep $t$ is injected through an adaptive LayerNorm [67] and the attention operator is augmented with 1D-RoPE [77] to encode positional information. To condition on text prompts, we encode them using the T5 [70] encoder and insert a cross attention layer attending to the text embeddings after each self attention operation [68]. The model contains 576M parameters and is trained to generate variable-length audio clips at 16kHz following previous work [32, 54]. We train it using the Flow Matching objective (Eq. (2)) over 100k iterations, with a learning rate of $43e\!-\!$ 4 and a batch size of 1024, distributed across 8 A100 GPUs. During inference, a pre-trained Vocoder [41] converts the generated Mel-spectrograms into waveforms.
Video Model. Given an RGB video $\mathbf{V}_{1}$, we flatten it to shape $\mathbb{R}^{T_{v}\times 3}$, where $T_{v}$ indicates the total number of video pixels. We adopt a DiT architecture symmetric to the audio model, and use 3D-RoPE [18]. Due to the computational costs associated with large scale video generators [99, 68] and the orthogonality between improvement of single-modality generation and AV-Link, we adopt a relatively small model (576M parameters) for our experiments, enabling efficient training and inference on the video modality. We found that when using a latent diffusion formulation, its motion and structural quality do not match the ones from an RGB model of the same size. Consequently, we opt to train a base video model to generate 5.16-second videos in the RGB space at a resolution of 36 x 64 pixels and 6 frames-per-second (FPS). The resolution is later increased to 512 x 288 at 24 FPS through a latent video upsampler operating in the latent space of a MAGVIT-v2 [102] autoencoder that is directly finetuned from the video model. The base video model is trained for 250k iterations with a batch size of 512, while the upsampler uses a batch size of 384 for 25k iterations. Both models are trained with a learning rate of $43e\!-\!$ 4 on 16 A100 GPUs.
@@ 3.3 Multimodal Fusion Block
Cross-modal generation requires a high-quality representation for the conditioning modality [57, 104]. In the context of V2A, obtaining high-quality video features has been the focus of recent work, which explored the usage of contrastive representations such as ImageBind [92], CLIP [104], MetaCLIP [68], or built ad hoc video representations [57, 104]. While such representations capture the overall video semantic, we observe that they lack precise audio-video temporal alignment, producing characteristic temporal misalignment (see Sec. 4.4). On the other hand, the video generator is capable of generating videos from scratch, implying that its activations contain semantically and temporally aligned information that can be leveraged for V2A. We thus propose a fully symmetric framework for V2A and A2V that leverages activations from the frozen generator of the conditioning modality to achieve aligned cross-modal generation using no ad-hoc feature extractors.
Fusion Blocks We address V2A and A2V by linking frozen audio and video generators $\mathcal{G}_{A}$, $\mathcal{G}_{V}$ through the Fusion Block depicted in Fig. 2. Consider the output activations of audio and video DiT blocks $\boldsymbol{x}_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$, $\boldsymbol{x}_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$ at some diffusion timestep $t_{a}$, $t_{v}$. We derive the activation of the conditional modality from a ground truth input (e.g. video for the V2A task). Initially, we project activations into a common dimension $D$ to produce $\bar{\boldsymbol{x}}_{a}\in\mathbb{R}^{T_{a}\times D}$, $\bar{\boldsymbol{x}}_{v}\in\mathbb{R}^{T_{v}\times D}$ respectively. A multi-head self attention operation is then applied to the concatenation of audio and video features $\{\bar{\boldsymbol{x}}_{a},\bar{\boldsymbol{x}}_{v}\}$. Finally, an MLP projects the output back to the original dimensionality, producing video-aware audio activations $\boldsymbol{\hat{x}}_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ and audio-aware video activations $\boldsymbol{\hat{x}}_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$ as:
@@
The Fusion Block presents a simple and symmetric design that equally treats the different modalities to condition the other modality.
Symmetric Feature Reinjection While the Fusion blocks effectively transfer the conditioning signal to the generated modality, we hypothesize that this signal can benefit from a continuous refinement driven by the generated modality. The symmetric architecture of the Fusion Blocks enables bidirectional information flow between the conditioning and generated modalities, rather than a one-way flow from the conditioning to the generated modality. Therefore, we propose to pass the conditional-aware (i.e. $\boldsymbol{\hat{x}}_{a}$ in the V2A task) and generation-aware activations (i.e. $\boldsymbol{\hat{x}}_{v}$ in V2A task) to subsequent DiT blocks within the respective frozen backbones, thereby injecting cross-modal information. In contrast to using static feature extractors [68, 57, 104], our design, which we call symmetric feature reinjection, dynamically improves the quality of the activations extracted from successive blocks of the conditioning modality.
Multimodal Temporal Alignment To facilitate temporal alignment between modalities, we use temporally-aligned 1D RoPE [77] embeddings in the self attention operation. We express 1D temporal RoPE as:
@@
Where $\eta_{a}$, $\eta_{v}$ are the number of tokens to represent one second of audio and video respectively. Intuitively, this approach rotates corresponding audio and video tokens by the same angle proportional to their temporal position, thus establishing temporal correspondence that facilitates temporally-synced multimodal information exchange. Crucially, this design aligns the modalities within the inserted Fusion Blocks while keeping the pretrained generators frozen. We show in Sec. 4.4 that this design is favored over previous temporal alignment approaches.
Time-Aware Feature Fusion We inject the information on the flow timesteps for audio $t_{a}$ and video $t_{v}$ through adaptive adaLN-Zero [66], with audio and video parameter sets receiving information on both timesteps. Such design creates conditioning features that are most suited for the current flow time for the generated modality, increasing the effectiveness of symmetric feature reinjection.
@@
Inference In the V2A setting, we consider the input conditioning video $\mathbf{V}_{1}$ and compute partially noised video $\mathbf{V}_{t_{v}}$ at the chosen flow time $t_{v}$, which we set to a fixed level during the entire inference phase to serve as input for the video backbone. As we later show in 4.4, it is crucial to inject some noise into the conditioning modality to extract useful features from the conditioning model. We sample $\mathbf{A}_{0}$ from Gaussian noise at $t_{a}\!=\!0$ and use velocity predictions from $\mathcal{G}_{V2A}$ to progressively denoise it into $\mathbf{A}_{1}$. We perform sampling using a first-order Euler solver with 64 sampling steps. Text prompts can optionally be used for both modalities to condition generation. We perform A2V inference completely symmetrically.